A home decor center was given a drawing on which a rectangular room measured 3 1/2 inches by 5 3/4 inches. The scale of the drawing was 1 inch respresented 5 feet. Find the dimensions of the room.
Answer by doukungfoo(195)

The scale we are given tells us that 1 inch = 5 feet.
So to find the feet that 3 1/2 inches represents just mulitply 3 1/2 by 5
And to find the feet that 5 3/4 inches represents just multiply 5 3/4 by 5
Ok so lets do the 3 1/2 first
3 1/2 * 5
convert 3 1/2 to an improper fraction
to do this multiply the denominator 2 times the whole number 3 and add the numerator 1. This is the numerator for the improper fraction. The denominator remains the same so
3 1/2 = 7/2
now we are ready to multiply by 5
7/2 * 5/1 = 35/2
convert back to a mixed number and you get 17 1/2
this is the first dimension: 17 1/2 feet
now do the same thing with 5 3/4
5 3/4 * 5
convert 5 3/4 to an improper fraction
to do this multiply the denominator 4 times the whole number 5 and add the numerator 3. This is the numerator for the improper fraction. The denominator remains the same so
5 3/4 = 23/4
now we are ready to multiply by 5
23/4 * 5/1 = 115/4
convert back to a mixed number and you get 28 3/4
this is the second dimension: 28 3/4 feet
So the actual dimensions of the room are 17 1/2 feet by 28 3/4 feet